# Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registrations

### v1

@@ Abstract
In this paper, we discuss Mahalanobis k-NN: a statistical lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds, either in the source or target point cloud. We tackle this by adopting Mahalanobis k-NN’s inherent property to capture the distribution of the local neighborhood and surficial geometry. Our method can be seamlessly integrated into any local-graph-based point cloud analysis method. In this paper, we focus on two distinct methodologies: Deep Closest Point (DCP) and Deep Universal Manifold Embedding (DeepUME). Our extensive benchmarking on the ModelNet40 and Faust datasets highlights the efficacy of the proposed method in point cloud registration tasks. Moreover, we establish for the first time that the features acquired through point cloud registration inherently can possess discriminative capabilities. This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task benchmarked on ModelNet40 and ScanObjectNN. The code is publicly available at https://github.com/TejasAnvekar/Mahalanobis-k-NN
Keywords:
Point Cloud Registration Metric Distance 3D Vision.
Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration task involving a target point cloud with only half the points compared to the source point cloud. In the visualization, highlighted boxes in red illustrate the limited performance of DCP-v1 [DCP] and DeepUME [DeepUME], while the highlighted boxes in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1 [DCP] and DeepUME [DeepUME].
@@ 1 Introduction
With the rapid progress of high-precision sensors like LiDAR [lidar] and Kinect [kinect], point clouds have become the prevalent data format for 3D representation. Due to sensor limitations in capturing only partial views, registration algorithms are crucial for amalgamating scans into comprehensive 3D scenes. Point cloud registration involves determining transformation matrices between point cloud pairs, facilitating the fusion of partial scans to produce coherent 3D representations.
@@
Registration is a fundamental task spanning various computational fields such as medical imaging [medicalImaging], robotics [robotics], autonomous driving [autoDriving], and computational chemistry [ComChe]. Estimating transformation matrices serves applications like 3D reconstruction [reconstruction] and precise 3D localization [Localization]. It underpins the development of intricate 3D maps for autonomous driving [autoDriving], environment reconstruction in robotics [robotics], and improved safety in mining. Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [autoDriving], ensuring precise positioning and interaction within the 3D environment.
@@
The registration of point clouds has garnered significant attention through closed-form [ICP, goICP, FGR, aligningPointCloudView, ume] and learning-based [pointnetLK, DCP, DeepGMR, DeepUME, rgm] methodologies. DCP [DCP] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [dgcnn] (Dynamic Graph Convolutional Neural Network). Subsequent steps involve an attention-based module and differentiable SVD modules for point-to-point registration. Conversely, DeepUME [DeepUME] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations. This space is leveraged to compute invariant per-point features that address real-world data sampling issues. Employing two weight-shared transformers and DGCNN [dgcnn], correspondences are established and fed into UME [ume] for rigid transformation parameter estimation. A limitation of both DCP [DCP] and DeepUME [DeepUME] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors. This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [dgcnn], which operates on graphs constructed using k-nearest neighbors (k-NN).
@@
We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate this issue. The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [MD1]. Our proposed Mahalanobis k-NN can be used as a plugin for any point-cloud registration method. In this paper, we propose mahalnobis versions of DCP-(v1,v2) [DCP] and DeepUME [DeepUME]. Both methods operate on diverse point-cloud representation spaces. This augmentation facilitates pointcloud registration as shown in Figure 1; across various publicly available 3D datasets, encompassing diverse benchmarking scenarios, including: 1) unseen-category evaluation, 2) robustness to various noise types, and 3) efficiency towards varying point densities. Additionally, we demonstrate the discriminative capabilities of the proposed Mahalanobis lens through point cloud few-shot classification tasks. In this context, models pre-trained for registration tasks leverage DGCNN [dgcnn] features for few-shot evaluation on both ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets. Finally, we summarize our contributions as follows:
@@
    • Deep Closest Point (DCP) [DCP] that operates on Euclidean coordinate space.
    • Deep Universal Manifold Embedding (DeepUME) [DeepUME] that operates on $\mathcal{SO}(3)$ invarient space.
@@
  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets through features extracted by DCP [DCP], we report up to an average of 20% increment when Mahalanobis is incorporated.
  • We will release our code to facilitate reproducibility and future research upon acceptance.
@@ 2 Related Works
Optimization Based. Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [comprehensiveEvalFeatureDescriptor, spinImages, aligningPointCloudView, RepeatabilityRobust3Dmatching, fastRobustDescriptor, rotationalContour]. Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [3Dsift] and the 3D Harris key-point detector [harris3D]. However, key-point matching in 3D presents challenges due to the absence of a regular sampling grid, artifacts, and sampling noise, leading to high outlier rates and localization errors. To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [ransac] and is refined using local optimization algorithms [ICP, scanRegistration3Dndt, registrationWithoutICP, iterativePointMatching]. Notably, DGR [DGR] follows a similar paradigm, but it incorporates learnable inlier detection. Researchers have proposed numerous works for handling outliers and noise [trimmedICP], formulating robust minimizers [robustRegistration2D3D], and devising more suitable distance metrics. The widely used Iterative Closest Point algorithm (ICP) [ICP, iterativePointMatching], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step. Over time, various ICP variants [efficientICP, goICP, sparseICP] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
@@
Learning Based. Registration approaches leverage learning-based techniques alongside closed-form methods. Pioneered by PointNet [pointnet] and advanced by DGCNN [dgcnn], learning from data for point cloud registration has become powerful (e.g., PointNet++, DCP [DCP], PRNet [prnet]). PointNetLK [pointnetLK] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm, while DCP uses attention and differentiable SVD for point-to-point registration. Recent work [rgm] introduces graph transformation and deep graph matching. Moreover, DeepUME [DeepUME] employ $\mathcal{SO}(3)$ invariant coordinate systems for global registration. DeepGMR [DeepGMR] tackles pose-invariant correspondences between raw point clouds and Gaussian-mixture models but suffers from noise sensitivity. DeepUME [DeepUME] fuses closed-form Universal Manifold Embedding (UME) and deep learning to handle sparse and unevenly sampled point clouds with large transformations. This framework is trained end-to-end and learns joint-resampling strategies and invariant features.
@@
DCP [DCP] and DeepUME [DeepUME] achieve impressive performance but struggle when the source or target point cloud densities are unequal. Both rely on DGCNN [dgcnn] for local feature-based correspondence estimation. However, DGCNN’s Euclidean metric for k-nearest neighbors lacks surface awareness. We propose Mahalanobis versions of DCP and DeepUME, introducing the Mahalanobis distance for surface-aware features and enhanced registration in real-world scenarios. Furthermore, our proposed approach demonstrates remarkable discriminative capabilities, as evident from its performance in point cloud few-shot classification tasks.
@@ 3 Problem Setting
The term $d\big{(}X,g(Y)\big{)}$ is a measure of the projection error between $X$ and the transformed $Y$, expressed as $d\big{(}X,RY+t\big{)}=\sum_{s=1}^{S}\left\|x_{s}-\big{(}Ry_{s}+t\big{)}\right\|_{2}$, where $s$ iterates over the $S$ pairs of corresponding points. The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [correspondance1, correspondance2]; conversely, accurate correspondences can be established given the optimal transformation matrix. However, the simultaneous resolution of both aspects poses a non-trivial challenge.
@@ 4 Method
In this paper, we focus on point cloud registration using DCP [DCP] and DeepUME [DeepUME] for global one-shot rigid transformations (contrasting iterative methods like ICP [ICP]). Both rely on DGCNN [dgcnn] for local features but use static k-nearest neighbor (k-NN) graphs, limiting semantic space creation. Unlike, DGCNN’s dynamic k-NN graphs, these methods restrict feature projection depth. We propose using the Mahalanobis distance [MD] to enhance k-NN, enabling surface-aware neighbor selection based on principal components. This extracts surface and corner features crucial for accurate matching and robust transformation estimation.
@@ 4.1 Mahalanobis k-NN
Figure 2: The illustration contrasts Euclidean distance fields (red) capturing spatial neighbors with Mahalanobis distance fields (blue) considering the underlying data distribution. A chair point cloud exemplifies this. The black query point is surrounded by Euclidean neighbors (red points) and Mahalanobis neighbors (blue points). The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surface points per the data distribution vital for precise feature matching.
In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [DCP] and DeepUME [DeepUME]. We modify DGCNN to construct a graph $\mathcal{G}^{\dagger}$ using Mahalanobis k-nearest neighbors (k-NN) for enhanced feature selection. DGCNN constructs a graph $\mathcal{G}$, applies nonlinearity for edge values, and performs vertex-wise aggregation ($\max$ or $\sum$) in each layer. Let $x_{i}^{l}$ be the embedding of point $i$ in the $l$-th layer, and $h_{\theta}^{l}$ be a nonlinear function parameterized by a shared multi-layer perceptron (MLP). DGCNN can be formulated as:
@@
We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation.2 to construct graph $\mathcal{G}^{\dagger}$. This graph prioritizes surface awareness through selection of surficial neighbors $\mathcal{N}_{i}^{\dagger}$ based on Mahalanobis distance as shown in Figure 2.
@@
The impact of Mahalanobis distance is visually depicted in Figure 2. This surficial awareness proves crucial in feature matching. Incorporating Mahalanobis distance into DCP  [DCP] and DeepUME  [DeepUME] yields superior performance in point cloud registration. Discriminative power is affirmed in point-cloud few-shot classification, with pre-trained DGCNN models from vanilla DCP and Mahalanobis DCP showcasing remarkable classification capabilities.
@@ 5 Experiments
In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [ICP], GO-ICP [goICP], FGR [FGR], PointNetLK [pointnetLK], DCP [DCP], and DeepUME [DeepUME] on publicly available datasets. Our evaluation encompasses a spectrum of benchmarking strategies, encompassing scenarios where one point cloud is notably sparser than the other. For a fair comparison, we implemented a vectorized Floyd-Warshall algorithm in PyTorch to compute geodesic distances and compare it to our proposed method for surface feature extraction. We further validate the Mahalanobis metric’s discriminative power in few-shot learning tasks and showcase robustness through endurance tests. Our model was trained using an Nvidia RTX 3090 GPU and PyTorch 1.11.

### v2

@@ Abstract
In this paper, we discuss Mahalanobis k-NN: A Statistical Lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds. We tackle this by adopting Mahalanobis k-NN’s inherent property to capture the distribution of the local neighborhood and surficial geometry. Our method can be seamlessly integrated into any local-graph-based point cloud analysis method. In this paper, we focus on two distinct methodologies: Deep Closest Point (DCP) and Deep Universal Manifold Embedding (DeepUME). Our extensive benchmarking on the ModelNet40 and FAUST datasets highlights the efficacy of the proposed method in point cloud registration tasks. Moreover, we establish for the first time that the features acquired through point cloud registration inherently can possess discriminative capabilities. This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task, benchmarked on ModelNet40 and ScanObjectNN.
Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration, involving a target point cloud with only half the points compared to the source point cloud. In the visualization, regions highlighted in red illustrate the limited performance of DCP-v1[42] and DeepUME[27], while the regions highlighted in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1[42] and DeepUME[27].
@@ 1 Introduction
With the rapid progress of high-precision sensors like LiDAR [13] and Kinect [52], point clouds are prevalent data format for 3D representation. Due to sensor limitations in capturing only partial views, registration algorithms are crucial for amalgamating scans into comprehensive 3D scenes. Point cloud registration involves determining transformation matrices between point cloud pairs, facilitating the fusion of partial scans to produce coherent 3D representations.
@@
Registration is a fundamental task spanning various computational fields such as medical imaging [9], robotics [32], autonomous driving [12], and computational chemistry [25]. Estimating transformation matrices serves applications like 3D reconstruction [21] and precise 3D localization [22]. It underpins the development of intricate 3D maps for autonomous driving [12], environment reconstruction in robotics [32], and improved safety in mining. Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [12], ensuring precise positioning and interaction within the 3D environment.
@@
The registration of point clouds has garnered significant attention through closed-form [4, 47, 53, 37, 15] and learning-based [2, 42, 50, 27, 18] methodologies. DCP [42] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [44] (Dynamic Graph Convolutional Neural Network). Subsequent steps involve an attention-based module and differentiable SVD modules for point-to-point registration. Conversely, DeepUME [27] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations. This space is leveraged to compute invariant per-point features that address real-world data sampling issues. Employing two weight-shared transformers and DGCNN [44], correspondences are established and fed into UME [15] for rigid transformation parameter estimation. A limitation of both DCP [42] and DeepUME [27] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors. This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [44], which operates on graphs constructed using k-nearest neighbors (k-NN).
@@
We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate aforementioned issue. The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [24]. Our proposed Mahalanobis k-NN can be used as a plugin for any point-cloud registration method. In this paper, we propose mahalnobis versions of DCP-(v1,v2) [42] and DeepUME [27]. Both methods operate on diverse point-cloud representation spaces. This augmentation facilitates pointcloud registration as shown in Figure 1; across various publicly available 3D datasets, encompassing diverse benchmarking scenarios, including: 1) unseen-category evaluation, 2) robustness to various noise types, and 3) efficiency towards varying point densities. Additionally, we demonstrate the discriminative capabilities of the proposed Mahalanobis lens through point cloud few-shot classification tasks. In this context, models pre-trained for registration tasks leverage DGCNN [44] features for few-shot evaluation on both ModelNet40 [45] and ScanObjectNN [41] datasets. Finally, we summarize our contributions as follows:
@@
    • Deep Closest Point (DCP) [42] that operates on Euclidean coordinate space.
    • Deep Universal Manifold Embedding (DeepUME) [27] that operates on $\mathcal{SO}(3)$ invariant space.
@@
  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [45] and ScanObjectNN [41] datasets through features extracted by DCP [42], we report up to an average of 20% increment when Mahalanobis is incorporated.
@@ 2 Related Works
Optimization Based. Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [19, 26, 37, 48, 46, 49]. Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [29] and the 3D Harris key-point detector [40]. However, key-point matching in 3D presents challenges due to the absence of a regular sampling grid, artifacts, and sampling noise, leading to high outlier rates and localization errors. To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [16] and is refined using local optimization algorithms [4, 30, 33, 51]. Notably, DGR [11] follows a similar paradigm, but it incorporates learnable inlier detection. Researchers have proposed numerous works for handling outliers and noise [10], formulating robust minimizers [17], and devising more suitable distance metrics. The widely used Iterative Closest Point algorithm (ICP) [4, 51], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step. Over time, various ICP variants [36, 47, 8] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
@@
Learning Based. Apart from optimization-based methods, registration approaches have also leveraged learning-based techniques. Pioneered by PointNet  [34] and further developed by DGCNN [44], learning from data in a task-specific manner has become a powerful tool in point cloud registration. These learned point cloud representations can be exploited for robust registration, as evidenced in studies like [35, 42, 43]. For instance, PointNetLK  [2] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm  [28], while DCP [42] addresses feature matching through attention-based modules and differentiable SVD modules for point-to-point registration. Recently, [18] introduces a transformation of point clouds into graphs and utilizes deep graph matching to extract deep feature soft correspondence matrices. To overcome this limitation, researchers have proposed approaches [27] employing $\mathcal{SO}(3)$ invariant coordinate systems to provide a global registration solution. Another method, DeepGMR  [50], deals with pose-invariant correspondences between raw point clouds and Gaussian mixture model (GMM) parameters, recovering transformations from the matched mixtures, although it exhibits performance degradation in the presence of sampling noise. Moreover, DeepUME [27] lies in its innovative fusion of the closed-form Universal Manifold Embedding (UME) method and a deep neural network to address the challenge of registering sparsely and unevenly sampled point clouds subjected to large transformations. This unified framework, trained end-to-end and without supervision, employs an $\mathcal{SO}(3)$-invariant coordinate system to learn joint-resampling strategies and invariant features.
@@
While DCP [42] and DeepUME [27] exhibit impressive performance across various real-world scenarios in point cloud registration, such as robustness to diverse noise types, they face limitations when dealing with imbalanced point cloud densities. Both methods rely on correspondence estimation based on local features using DGCNN [44]. However, DGCNN’s intrinsic lack of surface awareness due to its utilization of the Euclidean metric for k-nearest neighbors introduces challenges [38, 39]. To overcome this, authors in [38, 39] introduced using ellipsoid query but falls short as it requires extraneous querying. Counter to this we propose Mahalanobis versions of both DCP and DeepUME, introducing the Mahalanobis distance into k-NN to derive surface-aware features that enhance point cloud registration. Furthermore, our proposed approach demonstrates remarkable discriminative capabilities, as evident from its performance in point cloud few-shot classification tasks.
@@ 3 Problem Setting
The term $d\big{(}X,g(Y)\big{)}$ is a measure of the projection error between $X$ and the transformed $Y$, expressed as $d\big{(}X,RY+t\big{)}=\sum_{s=1}^{S}\left\|x_{s}-\big{(}Ry_{s}+t\big{)}\right\|_{2}$, where $s$ iterates over the $S$ pairs of corresponding points. The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [5, 6]; conversely, accurate correspondences can be established given the optimal transformation matrix. However, the simultaneous resolution of both aspects poses a non-trivial challenge.
@@ 4 Method
In this paper, we focus on two deep learning-based approaches, specifically DCP  [42] and DeepUME  [27], which address the aforementioned challenge as a global one-shot rigid transformation. This contrasts with the iterative nature of methods like ICP  [4]. Considering these methods, DGCNN  [44] is pivotal for extracting local features crucial for high-quality point matching. Both approaches modify the DGCNN framework to construct a static graph using k-nearest Neighbors (k-NN) for the point cloud. However, this limits the potential for complex semantic space creation. The conventional DGCNN recalculates the k-NN graph after each convolution, projecting features into higher-dimensional space, while the simplified k-NN graph imposition curtails semantic depth. To overcome this limitation, we propose harnessing the Mahalanobis distance [31] as a statistical lens for enhancing k-NN. This empowers querying based on principal components, emphasizing surface-level k-NN grouping. The Mahalanobis-based k-NN lens extracts features sensitive to surface and corner regions, vital for accurate feature matching and robust transformation estimation.
@@ 4.1 Mahalanobis k-NN
Figure 2: Illustration highlights the distinction between the Red gradients representing Euclidean distance fields that capture spatial neighbors and the Blue gradients depicting Mahalanobis distance fields, which consider neighbors concerning the underlying data distribution. To further emphasize the point, we present the influence of both Euclidean and Mahalanobis k-NN on a chair point cloud. The black query point is surrounded by Euclidean neighbors shown in red points and Mahalanobis neighbors in blue points. The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surficial points per the data distribution—vital for precise feature matching.
In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [42] and DeepUME [27]. We modify DGCNN to construct a graph $\mathcal{G}^{\dagger}$ using Mahalanobis k-nearest neighbors (k-NN) for enhanced feature selection. DGCNN constructs a graph $\mathcal{G}$, applies nonlinearity for edge values, and performs vertex-wise aggregation ($\max$ or $\sum$) in each layer. Let $x_{i}^{l}$ be the embedding of point $i$ in the $l$-th layer, and $h_{\theta}^{l}$ be a nonlinear function parameterized by a shared multi-layer perceptron (MLP). DGCNN can be formulated as:
@@
We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation 2 to construct graph $\mathcal{G}^{\dagger}$. This graph prioritizes surface awareness through selection of surficial neighbors $\mathcal{N}_{i}^{\dagger}$ based on Mahalanobis distance as shown in Figure 2.
@@
Advantages of Mahalanobis Distance. Unlike Euclidean distance which treats dimensions independently, Mahalanobis distance accounts for covariance, making it suitable for correlated feature datasets. It scales features invariantly, accommodating feature scale variations that affect Euclidean distances. This metric excels for datasets with non-spherical or elliptical clusters, considering data distribution shape. As data dimensionality increases, Mahalanobis distance is more resilient against the “curse of dimensionality,” particularly impactful on Euclidean distances. It demonstrates robustness to outliers, focusing on central data distribution. Moreover, Mahalanobis distance enables customized metric learning, adapting to dataset peculiarities for improved performance. In machine learning, it serves as a discriminative metric, enhancing class separation in classification. Its statistical interpretation enhances usability in data analysis. Moreover, it’s faster and more robust compared to geodesic-graph-based approaches.
The impact of Mahalanobis distance is visually depicted in Figure 2. This surficial awareness proves crucial in feature matching. Incorporating Mahalanobis distance into DCP  [42] and DeepUME  [27] yields superior performance in point cloud registration. Discriminative power is affirmed in point-cloud few-shot classification, with pre-trained DGCNN models from vanilla DCP and Mahalanobis DCP showcasing remarkable classification capabilities.
@@ 5 Experiments
In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [4], GO-ICP [47], FGR [53], PointNetLK [2], DCP [42], and DeepUME [27] on publicly available datasets. Our evaluation encompasses a spectrum of benchmarking strategies, encompassing scenarios where one point cloud is notably sparser than the other. For a fair comparison, we implemented a vectorized Floyd-Warshall algorithm in PyTorch to compute geodesic distances and compare it to our proposed method for surface feature extraction. We further validate the Mahalanobis metric’s discriminative power in few-shot learning tasks and showcase robustness through endurance tests. Our model was trained using an Nvidia RTX 3090 GPU and PyTorch 1.11.